# Pre-Training Representations of Binary Code Using Contrastive Learning

### v3

@@ Abstract
Binary code analysis and comprehension is critical to applications in reverse engineering and computer security tasks where source code is not available. Unfortunately, unlike source code, binary code lacks semantics and is more difficult for human engineers to understand and analyze. Limited work has explored incorporating multiple program representations. In this paper, we present ContraBin, a contrastive learning technique that integrates source code and comment information along with binaries to create an embedding capable of aiding binary analysis and comprehension tasks. Specifically, we present three components in ContraBin: (1) a primary contrastive learning method for initial pre-training, (2) a simplex interpolation method to integrate source code, comments, and binary code, and (3) an intermediate representation learning algorithm to train a binary code embedding. We evaluate the effectiveness of ContraBin through four indicative downstream tasks related to binary code: algorithmic functionality classification, function name recovery, code summarization, and reverse engineering. The results show that ContraBin considerably improves performance on all four tasks, measured by accuracy, mean of average precision, and BLEU scores as appropriate. ContraBin is the first language representation model to incorporate source code, binary code, and comments into contrastive code representation learning and is intended to contribute to the field of binary code analysis. The dataset used in this study is available for further research.
@@ 1 Introduction
Figure 2: Diagram of simplex interpolation and intermediate representation learning in ContraBin. It consists of four main components: (a) triplet data preparation, (b) simplex interpolation for gradual learning 2.2, (c) intermediate contrastive learning 2.3 and (d) task-specific fine tuning 2.5. After pre-training an embedding, the trainable encoder for binary code will be applied to the task-specific fine-tuning stage. In ContraBin, we use assembly code lifted from executable binary code. In (b), two encoded representations are randomly selected for interpolation to obtain an intermediate representation.
@@ 2.1 Primary Contrastive Learning
To mitigate this cold start problem, we adopt primary contrastive learning in the first step of model training. Specifically, we use three representations of a program: comments (i.e., documentation written by human experts), source code (i.e., as written by developers), and binary code (i.e., as emitted from a compiler), and randomly choose two of them to conduct contrastive learning, as shown in Figure 2. During this step, the simplex interpolation module is disabled: this approach is inspired by multi-view learning [18] in computer vision, in which an object can have $k$ views, and all of the views are exactly from the same object. In our model, either source code, binary code, or comment can represent the same program, so we treat the three representations (also denoted as modalities or views) of a program to enrich the information in binary code There are two steps in primary contrastive learning: manifold projection and batch-wise similarity comparison.
@@
Batch-wise similarity comparison we then adopt the loss function in CLIP [19] to update model by batch-wise similarity and dissimilarity between two program representations. Then, we compute the in-batch cross similarity between the anchored representation (either comment or source code) and trainable representations (binary code) using
$logits=H_{c/s}\otimes H_{b}^{T}$ (2)
where $logits$ denotes in-batch cross similarity between comment or source code representation $h_{c/s}$ and binary code representation $h_{b}$. In this paper, $\otimes$ denotes matrix multiplication. Then, we compute in-batch similarity of anchored and interpolated representations respectively by
${sim}_{c/s}=H_{c/s}\otimes H_{c/s}^{T},\quad{sim}_{b}=H_{b}\otimes H_{b}^{T},$ (3)
where ${sim}_{c/s}$ and ${sim}_{b}$ denote in-batch similarity of comment or source code and binary code. Next, we compute the softmax of the average of similarities between two representations by
$targets=({sim}_{c/s}+{sim}_{b})/{2},$ (4)
where $targets$ denotes the average between source code similarity ${sim}_{s}$ and binary code similarities ${sim}_{b}$. For the loss computation, we adopt Cross Entropy (CE) [20] to compute in-batch loss by
$\mathcal{L}=\textit{CE}(logits,targets),$ (5)
where $\mathcal{L}$ denotes the model loss in source code and binary code representations. This loss function can help to align the binary code embedding closer to the representations of comments and binary code in the manifold $\mathcal{M}$, thereby providing a well-trained binary code embedding as the starting point for subsequent steps.
@@
2.2 Simplex Interpolation for Gradual Learning
@@
After completing primary contrastive learning, we use simplex interpolation for gradual learning among the three representations, as shown in Figure 2 (a). Simplex interpolation is inspired by human translation (e.g., from source code to comment or vice versa) and the code compilation (or decompilation) process, in which conversion from one representation to the other involves intermediate thinking or analysis. While we acknowledge that the manifold assumption is a common tool in understanding deep learning phenomena, its application here serves primarily to align with our color boxes discussed from Section 2.1 to Section 2.4. This assumption provides a simplified framework to conceptualize the interpolation process, aiding in the explanation of our model’s behavior in binary code representation learning.
@@
Linear interpolation for direct learning We create two interpolation methods to generate inputs for contrastive learning: linear interpolation and non-linear interpolation, as shown in a unified Figure 3. Both interpolation methods follow the same interpolation function $\Gamma(a,b;\lambda)$ defined as
$\Gamma(A,B;\lambda)=\lambda\odot A+(1-\lambda)\odot B,$ (6)
where $A$ and $B$ are batch data representations, $\lambda\in[0,1]$ denotes the interpolation index, and $\odot$ indicates element-wise product. In ContraBin, we have $H_{l}=\Gamma(H_{1},H_{2};\lambda_{l})$ and $H_{n}=\Gamma(H_{1},H_{2};\lambda_{n})$ for linear and non-linear interpolation, respectively. Linear interpolation imitates the learning process between one thing to another, in which people make overall progress towards all knowledge points for a given task. In practice, we regard the interpolation index as a trainable parameter, learned by a neural network, denoted as
${\lambda}_{l}=f_{interp}^{l}(H_{1}\oplus H_{2}),$ (7)
where $f_{interp}^{l}$ denotes the interp NN function that learn the optimal interpolation index for both $H_{1}$ and $H_{2}$, and ${\lambda}_{l}$ denotes a linear interpolation index ranging from 0 to 1. Note that the Interp NN blocks in Figure 3 is just one of the possible implementations. Then, we apply linear interpolation onto the manifold vectors of source code and binary code, using
$H_{l}={\lambda}_{l}\odot H_{1}+(1-{\lambda}_{l})\odot H_{2},$ (8)
where $H_{l}$ denotes batch-wise intermediate representation generated by the linear interpolation. Here, we adopt the broadcast property in computation so the index number ${\lambda}_{l}$ will be expanded to the shape of $H_{1}$ and $H_{2}$ and become an item in vector multiplication.
Non-linear interpolation for customized learning We further consider a non-linear interpolation to capture complex semantics and enhance the model performance. Specifically, instead of regarding the interpolation index as a real number, we allow it to be a matrix with the identical shape of $H_{1}$ and $H_{2}$. In this case, we are able to capture the feature-wise and finer-grained information. Formally, ${\lambda}_{n}=f_{interp}^{n}(H_{1}\oplus H_{2})$ and $H_{n}={\lambda}_{n}\odot H_{1}+(1-{\lambda}_{n})\odot H_{2}$. Both interpolated embedding $H_{l}$ and $H_{n}$ serve as input to the intermediate contrastive learning.
2.3 Intermediate Contrastive Learning
@@
After obtaining the anchored and intermediate representation, we use an intermediate contrastive learning method to learn binary code embeddings. This is inspired by a continuation of Subsection 2.2, in which we obtain intermediate representations that contains an alternative view of the semantics of the same program. We treat these intermediate representations as unseen but reasonable information about the program.
Learn smartly The last component of ContraBin facilitates the learning process for analogy and extrapolation. It helps to decide the optimal way to combine knowledge.
As shown in Figure 4, we use each interpolated program representation $h_{i}$ (interpolation between source code projection $h_{s}$ and comment projection $h_{c}$ in the Figure) to compare with the other program representation (binary code projection $h_{b}$ in the Figure) in manifold $\mathcal{M}$. For simplicity, $h^{\prime}$ denotes all dissimilar projections (including intermediate and the other representations not from the same program in a batch).
We adopt a similar training process in Subsection 2.1, replacing two of the representations with intermediate and anchored representations to conduct intermediate contrastive learning.
2.4 Gradual Learning
Following the gradual learning literature, models learn most effectively when they follow a natural, progressive order as humans do [22, 23]. Specifically, models learn high-level data abstracts initially, then progress to more complicated embeddings. Following this intuition, we train our model via primary contrastive learning during cold start. After several epochs (10 epochs in our implementation), we switch the training pattern to contrastive learning based on linear simplex interpolation and non-linear simplex interpolation to further improve the generalizability of our model and finalize the training process, shown in Algorithm 1.
By integrating all three parts, ContraBin follows a natural learning curve (i.e., from simple to complex) to better understand the meaning of binary code.
@@ 2.5 Task-Specific Fine-tuning
For evaluating the performance of our pre-trained binary code embedding, we select four downstream tasks: (1) algorithmic functionality classification from binaries, (2) binary code function name recover, (3) binary code summarization, and (4) binary reverse engineering. We choose these four to cover two binary code analysis tasks (1 and 2) and two binary code comprehension tasks (3 and 4). Detailed descriptions are as follows:
Algorithmic functionality classification from binaries For this task, we classify a binary code input according to its functionality (i.e., “quicksort” is different from “md5 hash”). Categorizing the functionality of binary code blocks enables developers to isolate code blocks of interest, which improves readability, saves development time, and aids in reverse engineering [24]. Here, we classify a given binary into a specific category (e.g., sorting vs. hasing vs. search). We hypothesize that our approach will enable accurate functionality classification because we incorporate information from multiple representations of many programs.
Function name recovery In this task, we recover function names from sequences of binary code. In legacy systems, malware analysis, or in cases where source is unavailable, researchers and engineers must work with binary code directly or use tools to analyze binaries. However, stripped binary executables lack meaningful names and debugging information. Hence, automatic classification of the function names in binary code can substantially reduce the manual effort required for binary code analysis.
Code summarization In this task, we aim to translate binary code into an English summary. Summarizing the functionality of binary code can improve the comprehension of software and the effectiveness of software analyses, while offering the potential to carry out important security and vulnerability assessment, reverse engineering, and software maintenance tasks.
Reverse Engeering In this task, we aim to decompile binary code back into source code, providing the ability to conduct further analysis. While decompilation is currently supported by tools like Hex-Rays and Ghidra, decompiled code typically lacks valuable symbol names and differs substantially from canonical code written by developers. Transforming binaries to canoncical source code plays a crucial role in comprehending binary code as it facilitates the examination of previously created software. By reversing the code, researchers gain insight into how it operates, identify problems, and make improvements to its functionality. This reverse engineering of binary to source facilittes security assessments, defect localization, and human comprehension of the original binary code. We categorize these tasks in Table I.
TABLE I: Categorization of four downstream tasks.
Binary Code Analysis Tasks Task Object
Algorithmic functionality classification (1) Analysis
Function name recovery (2) Analysis
Code summarization (3) Comprehension
Reverse engineering (4) Comprehension

### v4

@@ Abstract
Binary code analysis and comprehension is critical to applications in reverse engineering and computer security tasks where source code is not available. Unfortunately, unlike source code, binary code lacks semantics and is more difficult for human engineers to understand and analyze. In this paper, we present ContraBin, a contrastive learning technique that integrates source code and comment information along with binaries to create an embedding capable of aiding binary analysis and comprehension tasks. Specifically, we present three components in ContraBin: (1) a primary contrastive learning method for initial pre-training, (2) a simplex interpolation method to integrate source code, comments, and binary code, and (3) an intermediate representation learning algorithm to train a binary code embedding. We further analyze the impact of human-written and synthetic comments on binary code comprehension tasks, revealing a significant performance disparity. While synthetic comments provide substantial benefits, human-written comments are found to introduce noise, even resulting in performance drops compared to using no comments. These findings reshape the narrative around the role of comment types in binary code analysis. We evaluate the effectiveness of ContraBin through four indicative downstream tasks related to binary code: algorithmic functionality classification, function name recovery, code summarization, and reverse engineering. The results show that ContraBin considerably improves performance on all four tasks, measured by accuracy, mean of average precision, and BLEU scores as appropriate. ContraBin is the first language representation model to incorporate source code, binary code, and comments into contrastive code representation learning and is intended to contribute to the field of binary code analysis. The dataset used in this study is available for further research.
@@ 1 Introduction
    We provide an in-depth analysis of the impact of human-written and synthetic comments on binary code comprehension tasks, demonstrating that synthetic comments significantly enhance performance, while human-written comments can introduce noise and reduce model effectiveness compared to using no comments.
  • $\bullet$
@@
Figure 2: Overview of ContraBin’s Training Framework. The framework consists of four components: (a) Triplet Data Preparation, where triplet data (source code, binary code, and comments) are created through compilation and machine-generated comments; (b) and (c) are collectively structured into two learning stages: (1) Primary Contrastive Learning, which selects a single representation (naive SI) and applies batch-wise similarity and dissimilarity; and (2) Secondary Contrastive Learning, which generates interpolated representations using simplex interpolation (SI) and further aligns them. Finally, (d) Task-Specific Fine-Tuning applies the pre-trained embeddings to downstream tasks.
@@ 2.1 Primary Contrastive Learning
To mitigate this cold start problem, we adopt primary contrastive learning in the first step of model training. Specifically, we use three representations of a program: comments (i.e., summarization generated by code models), source code (i.e., as written by developers), and binary code (i.e., as emitted from a compiler), and randomly choose two of them to conduct contrastive learning, as shown in Figure 2. During this step, the simplex interpolation module is disabled: this approach is inspired by multi-view learning [18] in computer vision, in which an object can have $k$ views, and all of the views are exactly from the same object. In our model, either source code, binary code, or comment can represent the same program, so we treat the three representations (also denoted as modalities or views) of a program to enrich the information in binary code There are two steps in primary contrastive learning: manifold projection and batch-wise similarity comparison.
@@
Batch-wise similarity comparison To align binary code representations with their corresponding source code and comments, we adopt a loss function inspired by CLIP [19]. This method leverages batch-wise similarity and dissimilarity between different program representations to update the model.
Specifically, we compute the in-batch cross-similarity between the anchored representation (either comment or source code) and the trainable binary code representation as follows:
$\text{logits}=H_{c/s}H_{b}^{T},$ (2)
where $H_{c/s}$ and $H_{b}$ represent the embeddings of comments or source code and binary code, respectively, and the notation $H_{c/s}H_{b}^{T}$ indicates standard matrix multiplication. The resulting logits quantify the in-batch cross-similarity.
To ensure compatibility with the cross-entropy loss, we convert the logits into a probability distribution using the softmax function:
$P=\text{softmax}(\text{logits},\text{dim}=-1).$ (3)
Next, we compute the in-batch similarity of the representations within each modality:
$\text{sim}_{c/s}=H_{c/s}H_{c/s}^{T},\quad\text{sim}_{b}=H_{b}H_{b}^{T},$ (4)
where $\text{sim}_{c/s}$ and $\text{sim}_{b}$ represent the intra-batch similarities of the comment/source code and binary code embeddings, respectively.
To construct the target distribution for the loss computation, we take the average of the intra-batch similarities and apply the softmax function:
$\text{targets}=\text{softmax}\left(\frac{\text{sim}_{c/s}+\text{sim}_{b}}{2},\text{dim}=-1\right).$ (5)
The loss function is then computed using the cross-entropy (CE) between the probability distributions of $P$ and targets:
$\mathcal{L}=\textit{CE}(P,\text{targets}).$ (6)
By minimizing this loss, the model learns to align semantically similar binary code, source code, and comment embeddings in a unified embedding space. Minimizing the cross-entropy penalizes differences between $P$ and targets, forcing the two distributions to become more similar. As targets reflects the similarity structure of embeddings, this process aligns $P$ with it, effectively concentrating the distributions (reducing entropy) and pulling semantically related concepts closer in the embedding space.
@@
2.2 Secondary Contrastive Learning
To enhance the representation quality of binary code embeddings, secondary contrastive learning builds on primary contrastive learning by integrating simplex interpolation and intermediate contrastive learning. In primary contrastive learning, naive interpolation is used, where a single representation is selected without applying interpolation techniques (as illustrated in Figure 2). Secondary contrastive learning extends this by introducing two advanced simplex interpolation methods: linear interpolation for direct learning and non-linear interpolation for customized learning. These methods generate diverse and expressive views of program representations, enabling the model to better capture semantic relationships among source code, binary code, and comments. Intermediate contrastive learning then operates across all views generated by the simplex interpolation methods, providing a novel mechanism to align semantically related representations while preserving distinctions among unrelated ones. Figure 2 illustrates the integration and interaction of these components within a unified training pipeline.
@@ 2.2 Secondary Contrastive Learning
Simplex interpolation for gradual learning After completing primary contrastive learning, we use simplex interpolation for gradual learning among the three representations, as shown in Figure 2 (a). Simplex interpolation is inspired by human translation (e.g., from source code to comment or vice versa) and the code compilation (or decompilation) process, in which conversion from one representation to the other involves intermediate thinking or analysis. While we acknowledge that the manifold assumption is a common tool in understanding deep learning phenomena, its application here serves primarily to align with our color boxes discussed from Section 2.1 to Section 2.4. This assumption provides a simplified framework to conceptualize the interpolation process, aiding in the explanation of our model’s behavior in binary code representation learning.
@@
Linear interpolation for direct learning We create two interpolation methods to generate inputs for contrastive learning: linear interpolation and non-linear interpolation, as shown in a unified Figure 3. Both interpolation methods follow the same interpolation function $\Gamma(A,B;\lambda)$ defined as:
$\Gamma(A,B;\lambda)=\lambda\cdot A+(1-\lambda)\cdot B,$ (7)
where $A$ and $B$ are batch data representations, $\lambda\in[0,1]$ denotes the interpolation index, and $\cdot$ indicates element-wise multiplication. In ContraBin, we define $H_{l}=\Gamma(H_{1},H_{2};\lambda_{l})$ and $H_{n}=\Gamma(H_{1},H_{2};\lambda_{n})$ for linear and non-linear interpolation, respectively. Linear interpolation imitates the learning process between one thing to another, where the model progresses towards all knowledge points for a given task. In practice, we regard the interpolation index as a trainable parameter, learned by a neural network, denoted as:
$\lambda_{l}=f_{interp}^{l}(H_{1}+H_{2}),$ (8)
where $f_{interp}^{l}$ denotes the interpolation neural network function that learns the optimal interpolation index for $H_{1}$ and $H_{2}$, and $\lambda_{l}$ denotes a linear interpolation index ranging from 0 to 1. Note that the Interp NN blocks in Figure 3 are one of the possible implementations. We apply linear interpolation to the manifold vectors of source code and binary code using:
$H_{l}=\lambda_{l}\cdot H_{1}+(1-\lambda_{l})\cdot H_{2},$ (9)
where $H_{l}$ denotes the batch-wise intermediate representation generated by the linear interpolation. The broadcast property is adopted so that the scalar $\lambda_{l}$ expands to the shape of $H_{1}$ and $H_{2}$, enabling element-wise multiplication.
Non-linear interpolation for customized learning We further consider non-linear interpolation to capture complex semantics and enhance model performance. Unlike linear interpolation, which uses a scalar $\lambda$, non-linear interpolation treats the interpolation index as a matrix with the same shape as $H_{1}$ and $H_{2}$, allowing finer-grained and feature-wise control. Formally, this is defined as:
$\lambda_{n}=f_{interp}^{n}(H_{1}+H_{2}),$ (10)
$H_{n}=\lambda_{n}\cdot H_{1}+(1-\lambda_{n})\cdot H_{2},$ (11)
where $\lambda_{n}$ is a matrix representing the non-linear interpolation index and $f_{interp}^{n}$ is the corresponding neural network function. Both interpolated embeddings $H_{l}$ and $H_{n}$ serve as input to the intermediate contrastive learning step.
@@
Intermediate contrastive learning for representation refinement To further refine the binary code embeddings, we introduce intermediate contrastive learning, which leverages interpolated representations generated in Subsection 2.2. These interpolations provide an alternative view of the program’s semantics, acting as unseen yet reasonable information about the same program. This step builds on the concept of contrastive learning by incorporating these interpolated representations into the training process.
As illustrated in Figure 4, the interpolated representation $h_{i}$ (a combination of source code projection $h_{s}$ and comment projection $h_{c}$) is compared against the binary code projection $h_{b}$ in the embedding manifold $\mathcal{M}$. Similarly, $h^{\prime}$ denotes all dissimilar projections, which include other representations in the batch that do not belong to the same program.
To compute the intermediate contrastive loss, we define
$Z_{i}=\sum_{h^{\prime}\in\mathcal{B}^{\prime}}\exp(\text{sim}(h_{i},h^{\prime})/\tau),$
where $\mathcal{B}^{\prime}$ denotes the set of all dissimilar projections in the batch, $\text{sim}(\cdot,\cdot)$ represents the cosine similarity between two embeddings, and $\tau$ is a temperature parameter. The loss is then computed as:
$\mathcal{L}_{\text{intermediate}}=-\frac{1}{N}\sum_{i=1}^{N}\log\Bigg{(}\frac{\exp(\text{sim}(h_{i},h_{b})/\tau)}{\exp(\text{sim}(h_{i},h_{b})/\tau)+Z_{i}}\Bigg{)},$
where $N$ is the batch size. This formulation encourages the interpolated representation $h_{i}$ to align closely with the corresponding binary code projection $h_{b}$, while maintaining separation from unrelated projections $h^{\prime}$. By introducing $Z_{i}$, we explicitly represent the normalization factor in the denominator, which simplifies the interpretation of the loss and enhances its clarity.
Intermediate Representations in Context Intermediate contrastive learning refines embeddings by aligning related representations and separating unrelated ones, enabling ContraBin to better capture semantic relationships for improved learning.
To stabilize training, we adopt a stop-gradient mechanism for the anchored representation during optimization, detaching the gradient of $h_{s}$ or $h_{c}$ when used as the anchor to prevent instability caused by intermediate representations. This mechanism facilitates the integration of intermediate contrastive learning into the overall training process, ensuring that interpolated views contribute effectively to refining binary code embeddings. Building on this foundation, the subsequent gradual learning step further enhances the model’s ability to capture nuanced semantic relationships across diverse program representations.
Gradual learning Following the gradual learning literature, models learn most effectively when they follow a natural, progressive order as humans do [21, 22]. Specifically, models learn high-level data abstracts initially, then progress to more complicated embeddings. Following this intuition, we train our model via primary contrastive learning during cold start. After several epochs (10 epochs in our implementation), we switch the training pattern to contrastive learning based on linear simplex interpolation and non-linear simplex interpolation to further improve the generalizability of our model and finalize the training process, shown in Algorithm 1. By integrating all three parts, ContraBin follows a natural learning curve (i.e., from simple to complex) to better understand the meaning of binary code.
@@ 2.3 Task-Specific Fine-Tuning
To evaluate the effectiveness of our pre-trained binary code embeddings, we apply them to four downstream tasks: (1) algorithmic functionality classification from binaries, (2) binary code function name recovery, (3) binary code summarization, and (4) binary reverse engineering. These tasks comprehensively span two domains: binary code analysis (tasks 1 and 2) and binary code comprehension (tasks 3 and 4). By addressing a diverse set of challenges, our fine-tuning approach demonstrates the versatility and robustness of the learned embeddings.
Fine-tuning is performed using task-specific objectives, each tailored to the requirements of the respective downstream task. For classification tasks, including algorithmic functionality classification and function name recovery, we employ a cross-entropy loss function:
$\mathcal{L}_{\text{classification}}=-\frac{1}{N}\sum_{i=1}^{N}y_{i}\log\hat{y}_{i},$
where $y_{i}$ is the true label, $\hat{y}_{i}$ is the predicted probability, and $N$ is the batch size. This objective encourages the model to predict accurate labels for binary code representations.
For summarization and reverse engineering tasks, we treat the problem as a sequence-to-sequence generation task. Here, the loss function used is a token-level cross-entropy loss:
$\mathcal{L}_{\text{seq2seq}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}y_{i,t}\log\hat{y}_{i,t},$
where $T$ represents the sequence length, $y_{i,t}$ is the true token at time step $t$, and $\hat{y}_{i,t}$ is the predicted probability for that token. This formulation ensures the model learns to generate accurate textual summaries and source code representations.
All four tasks are adapted from the CodeXGlue [^3] https://github.com/microsoft/CodeXGLUE benchmark, with modifications for binary code analysis. For instance, the binary functionality classification and function name recovery tasks are based on their corresponding tasks in CodeXGlue but focus on pre-compiled binary representations rather than source code. Similarly, binary code summarization and reverse engineering extend the summarization and translation tasks in CodeXGlue to the domain of binary-to-natural language and binary-to-source code transformations, respectively. Detailed descriptions for each task are provided below.
TABLE I: Categorization and objectives of downstream tasks.
Task Domain Objective
Algorithmic functionality classification (1) Analysis Categorize binaries by functionality
Function name recovery (2) Analysis Predict meaningful function names
Code summarization (3) Comprehension Generate concise summaries for binaries
Reverse engineering (4) Comprehension Decompile binaries into source code
Algorithmic Functionality Classification from Binaries This task involves classifying binary code according to its functionality, such as distinguishing ”quicksort” from ”md5 hash.” Accurate classification assists developers in isolating specific code blocks, improving readability, saving development time, and aiding in reverse engineering [23]. For example, given a binary, the goal is to categorize it into classes like sorting, hashing, or searching. As shown in Table I, this task is categorized under ”Analysis” and represents a key aspect of binary code analysis, where our method leverages multiple program representations for robust classification.
Function Name Recovery Recovering function names from binary code sequences is crucial for stripped binaries that lack meaningful names and debugging information. This task, categorized as ”Analysis” in Table I, significantly reduces manual effort in reverse engineering and malware analysis scenarios, particularly when source code is unavailable. By predicting function names effectively, our embeddings restore clarity to binary code, enhancing the overall analysis workflow.
Code Summarization This task translates binary code into concise English summaries, aiding comprehension and facilitating security assessments, reverse engineering, and software maintenance. As outlined in Table I, code summarization falls under the ”Comprehension” category, emphasizing its role in improving understanding of binary functionality and enabling more effective software analysis through succinct summaries.
Reverse Engineering Decompiling binary code into source code is essential for tasks like security assessments, defect localization, and software comprehension. While tools like Hex-Rays and Ghidra offer decompilation support, they often generate code that lacks symbolic names and deviates from canonical developer-written formats. Reverse engineering, classified under ”Comprehension” in Table I, benefits from our embeddings by transforming binaries into more interpretable source code, providing valuable insights into the original binaries.
We summarize the categorization of these tasks, their objectives, and their respective focus areas in Table I, which groups the tasks into ”Analysis” and ”Comprehension” domains.
2.4 Parallelism Between the Presented Steps and the Learning Stages
To ensure a clear understanding of the connections between the training steps and the overall learning pipeline, we explicitly describe the parallelism between the training stages and learning steps in ContraBin. Each stage in our framework directly contributes to a specific aspect of the learning process, as illustrated in Figure 2:
  • Primary Contrastive Learning Step: This step corresponds to the initial stage of representation learning, where naive interpolation (selecting a single representation from source code or comments) is used. It serves as the foundational phase to align binary code with these modalities in the embedding space without introducing intermediate representations.
  • Secondary Contrastive Learning Step: This step builds upon the foundational alignment by introducing simplex interpolation, where representations of source code and comments are interpolated to generate intermediate views. These interpolated representations provide a nuanced semantic bridge, which, when combined with intermediate contrastive learning, allows the model to refine its embedding space. By focusing on relationships between source code and comments, this step ensures that binary code embeddings are enriched through contextually meaningful relationships.
  • Task-Specific Fine-Tuning Step: In this final stage, the pre-trained embeddings are applied to downstream tasks. This step translates the learned embeddings into practical utility, demonstrating their effectiveness across binary analysis and comprehension tasks.
The transitions between these steps ensure a seamless progression from foundational alignment to representation refinement and practical application. This structured pipeline integrates the learning stages cohesively, as depicted in Figure 2 and aligned with the corresponding text sections.